### v1

# Better Understandings and Configurations in MaxSAT Local Search Solvers via Anytime Performance Analysis

Abstract
Though numerous solvers have been proposed for the MaxSAT problem, and the benchmark environment such as MaxSAT Evaluations provides a platform for the comparison of the state-of-the-art solvers, existing assessments were usually evaluated based on the quality, e.g., fitness, of the best-found solutions obtained within a given running time budget. However, concerning solely the final obtained solutions regarding specific time budgets may restrict us from comprehending the behavior of the solvers along the convergence process. This paper demonstrates that Empirical Cumulative Distribution Functions can be used to compare MaxSAT local search solvers’ anytime performance across multiple problem instances and various time budgets. The assessment reveals distinctions in solvers’ performance and displays that the (dis)advantages of solvers adjust along different running times. This work also exhibits that the quantitative and high variance assessment of anytime performance can guide machines, i.e., automatic configurators, to search for better parameter settings. Our experimental results show that the hyperparameter optimization tool, i.e., SMAC, generally achieves better parameter settings of local search when using the anytime performance as the cost function, compared to using the fitness of the best-found solutions.
1 Introduction
The Maximum Satisfiability (MaxSAT) problem is the optimization version of the influential Boolean Satisfiability (SAT) problem, aiming at finding the maximum number of satisfied clauses. Partial MaxSAT, which is an important generalization of MaxSAT, divides the clauses into hard and soft ones, and a feasible solution requires all hard clauses to be satisfied. Among various complete and incomplete solvers developed for solving MaxSAT, local search (LS), which commonly follows an iterative framework, is a significant category of incomplete solvers. While new methods have been continuously proposed, the corresponding comparison results are presented in numerous documents. The famous MaxSAT Evaluations Berg et al. (2023) provide a fine platform for practitioners to understand the performance of state-of-the-art solvers, and the existing comparisons among MaxSAT solvers mainly concern the quality, i.e., fitness, of best solutions obtained within a given time budget, namely, cutoff time AlKasem and Menai (2021); Chu et al. (2023); Cai et al. (2014, 2017), which is a reasonable option when we require a concrete judgment for the competition of various solvers. However, the assessments concerning best-found solutions may produce bias in the cutoff time and hinder us from understanding the behavior of solvers during the iterative optimization process. Though some work Zheng et al. (2022) investigated the convergence process when analyzing LS solvers’ behavior, the current mechanism usually plots individual convergence lines for particular instances. Hickey and Beacchus’s work Hickey and Bacchus (2022) addressed the anytime performance of MaxSAT solvers, but their work focused on proposing a method that was evaluated based on the best-found solutions concerning multiple cutoff times without addressing the issue of measuring anytime performance.
Two perspectives, i.e., fixed-budget and fixed-target, are commonly considered when measuring iterative algorithms’ performance. The fixed-budget approach is usually applied to problems in which prior knowledge about the performance, e.g., the fitness scale, is unavailable, and it can assess the order of the obtained solutions. The fixed-target approach is commonly applied in well-understood practical scenarios or benchmarking environments, and it uses quantitative indicators to explain how fast an algorithm obtains a solution meeting specific requirements. However, we usually need to consider both perspectives when developing an algorithm for practical problems. As an alternative addressing concerns of the two mentioned approaches, an anytime performance metric, which examines algorithms’ performance based on empirical cumulative distribution functions (ECDF) of a set of cutoff times, has recently presented its superiority in black-box optimization benchmarking scenarios Hansen et al. (2016, 2022). The ECDF value at a given time is computed based on the fraction of solved problem instances or obtained solutions reaching the required target, e.g., fitness. In this way, we can obtain a ratio scaled value independent of the scale of solution fitness, and results regarding multiple cutoff times can be reasonably aggregated.
This paper introduces the first anytime performance assessments of MaxSAT LS solvers. Our anytime performance assessments reveal intriguing behaviors that were invisible for fixed-budget assessments. The observations can provide valuable insights for improving the design and parameter settings of future LS solvers. The hybrid solvers, which are the current state-of-the-art MaxSAT solvers, operate complete solvers and LS solvers independently and alternatively. Therefore, this work focusing on analyzing LS may contribute to potential enhancements of hybrid methods. In addition, the assessment techniques employed in this work can be transparent to both complete and hybrid solvers. In practice, our results reveal (dis)advantages of the solvers across different problem instances and demonstrate algorithms’ convergence progress aggregated across multiple instances. Apart from providing a comprehensive assessment that can inspire experts to improve the design of algorithms further, the quantitative assessment also creates an alternative way for algorithmic tools, e.g., hyperparameter optimizers, to recognize algorithms’ performance.
Since contemporary algorithms, including LS, are mostly parametric, hyperparameter optimization (HPO) is vital for robust and competitive performance. For example, the MaxSAT Evaluation competition submissions are usually embedded with fitting parameter settings. Meanwhile, automatic HPO tools have been commonly applied to achieve those fine-tuned settings. The HPO scenario for MaxSAT solvers usually addresses the algorithm configuration (AC) problem, which aims to find a configuration minimizing a cost function $e$. A configuration will be executed to evaluate the cost function $e$ until a cutoff time $\kappa$. Based on theoretical analysis of evolutionary algorithms, recent work Hall et al. (2022) has studied the impact of $\kappa$ on the performance of HPO, considering both scenarios using fixed-target and fixed-budget approaches as $e$. Tuning algorithms’ anytime performance has been addressed in the recent work Ye et al. (2022), which compared the HPO scenarios of tuning fixed-target and anytime performances. While the mentioned studies were tested on classic benchmark problems such as OneMax, López-Ibáñez and Stützle applied the hypervolume considering two objectives, i.e., the best-found solution quality and the corresponding cutoff time, to take into account the anytime behavior of algorithms López-Ibáñez and Stützle (2014), and the method was tested on a MAX-MIN Ant system for Traveling Salesperson Problems and SCIP Achterberg (2009). In this work, we investigate tuning the anytime performance of a MaxSAT LS solver, and the experimental results suggest that anytime performance is a better mechanism than the fixed-budget one for the cost function of HPO.
The main contributions of this work are as follows:
  • We provide the anytime performance assessment of four state-of-the-art MaxSAT LS solvers. We aggregate ECDF performance across multiple problem instances and compare the solvers’ performance concerning various cutoff times. The assessment illustrates that ECDF, as a universal technique with a ratio scale, can distinguish the solvers that are considered identical when comparing the fixed-budget performance. Moreover, we observe that the solvers’ performance varies along the optimization process while investigating their performance in particular instances, addressing the necessity of anytime performance assessment.
  • We suggest tuning anytime performance can obtain better parameter settings compared to the commonly applied fixed-budget performance tuning in the MaxSAT community. By tuning the ECDF of LS solvers, SMAC Lindauer et al. (2022) can obtain better configurations in terms of anytime performance (i.e., ECDFs) and fixed-budget performance (i.e., scores based on the best-found solution quality). Note that the mechanism of computing incremental aggregated ECDFs for HPO used in this paper can be applied to other scenarios without interfacing with the HPO framework.
2 Preliminaries
2.1 Problem Definition
Given a set of $n$ Boolean variables $V=\{x_{1},x_{2},\ldots,x_{n}\}$, a literal $l$ is either a variable $x$ or its negation $\neg x$, and a clause is a disjunction of literals, i.e., $c=l_{1}\lor l_{2}\ldots\lor l_{k}$, where $k$ is the length of $c$. Given an assignment $\alpha$ mapping Boolean values to each variable in $V$, a clause $c$ is satisfied if at least one literal in $c$ is True; otherwise, it is falsified. Given a conjunction normal form (CNF) $F=c_{1}\land c_{2}\land\ldots\land c_{m}$, where $m$ is the number of clauses, the MaxSAT problem is to find an assignment that maximizes the number of satisfied clauses. Partial MaxSAT (PMS) is a variant of MaxSAT that divides clauses into hard ones and soft ones, and a feasible assignment $\alpha$ requires all hard clauses to be satisfied. For the weighted PMS (WPMS) problem, each clause is associated with a weight $w(c)>0$, and the problem is to find a feasible $\alpha$ that maximizes the total weights of satisfied clauses (namely, minimizes the total weights of falsified clauses). In this paper, we denote (W)PMS as a constrained optimization problem minimizing the $cost(\alpha)$ that is the total weight of falsified clauses. Nowadays, MaxSAT solvers are commonly developed and tested for (W)PMS, and in this paper, we refer to MaxSAT as (W)PMS.
2.2 Algorithms
Local search (LS) is a popular category of incomplete algorithms for MaxSAT. The common procedure of LS samples an initial solution, i.e., assignment $\alpha$, and then follows an optimization loop, creating new assignments by flipping one or multiple variables $x$ in $\alpha$ iteratively until reaching the termination condition (e.g., using out the cutoff time). Variants of LS solvers usually work on the strategies of picking one or multiple variable(s) $x$ to be flipped at each iteration.
The classic SATLike solver Lei and Cai (2018) proposed a clause weighting scheme, which introduces a bias towards certain variables $x$ and determines which one(s) to be flipped. The newest extension of SATLike, SATLike3.0 Cai and Lei (2020), is one of the state-of-the-art LS solvers for MaxSAT and has been included in many recent works for comparisons AlKasem and Menai (2021). NuWLS Chu et al. (2023) proposed a new clause weighting scheme handling hard and soft clauses separately and worked on the weight initialization based on the framework of SATLike3.0. Differently from other methods that flip one variable at each iteration, MaxFPS Zheng et al. (2023) applies a farsighted probabilistic sampling method flipping a pair of variables. Moreover, instead of using the clause weighting technique, BandMax Zheng et al. (2022) introduced the method of multi-armed bandit to select clauses to be satisfied, correspondingly deciding which variable to be flipped.
In this paper, we select these four state-of-the-art methods, SATLike3.0, BandMax, NuWLS, and MaxFPS, for our first assessment of anytime performance of MaxSAT LS solvers.
3 Performance Assessment
We test the four state-of-the-art LS solvers on the 2022 and 2023 anytime tracks of MaxSAT Evaluations, including weighted (MSE23-w and MSE22-w) and unweighted MaxSAT (MSE23-uw and MSE22-uw). All the tested solvers are implemented in C++ and complied with g++ 9.4.0. All experiments are conducted in Ubuntu 20.04.4 with the AMD EPYC 7763 CPU and 1TB RAM. The cutoff time is set as $300s$ for each run following the suggestion of MaxSAT Evaluations, and we conduct a solver of ten independent runs for each instance. Additional detailed experimental results of the cutoff time $60s$, which is another suggestion of MaxSAT Evaluations, are available in our repository (see Section 6).
We introduce in the following the fixed-budget performance and anytime performance of the solvers. The former is commonly applied in the MaxSAT community, and the latter is first studied for MaxSAT in this work. The fixed-target performance is not addressed because MaxSAT is an NP-hard problem, and its optima are unknown.
3.1 Fixed-budget Performance
Fixed-budget performance metric.
Existing work on LS solvers for MaxSAT usually reports their experimental results based on (1) the number of winning instances (“#win”) and (2) the score based on the competing solution Berg et al. (2023); AlKasem and Menai (2021). The number of winning instances presents the scope in the tested instances where a solver obtains the best solution among all competing solvers, and the score denotes the ratio between the solution quality obtained by a solver and the best-competing solver:
$score(i)=\frac{\text{cost of the best solution for }i+1}{\text{cost of the found solution for }i+1}\in[0,1]$ (1)
$score(i)=0$ indicates that a solver can not find a feasible solution solving the instance $i$. These two metrics are both calculated based on the best-found solution within a cutoff time. Consequently, such assessment outcomes may deviate along different cutoff times. Note that we are not critical of using these measures, as they can provide concrete comparison conclusions for algorithm performance orders, as shown in amounts of published papers and previous MaxSAT Evaluation competitions.
Experimental Results.
We present in Table 1 the number of instances in which a solver obtains the best solution (that is achieved by all runs of the competing solvers) in one of the tested ten runs, followed by the number (in brackets) of instances in which a solver obtains the best average results of ten runs among the competing solvers. Results are categorized for the four tested instance sets. Due to the space issue, we denote SATLike as SATLike3.0 in the following text. According to Table 1, NuWLS obtains significant advantages against the other solvers, and BandMax and MaxFPS show similar performance across the four tested benchmark tracks. BandMax and MaxFPS win on more instances than SATLike when comparing the best of ten runs, but such advantages diminish when comparing the average of ten runs (as shown in the numbers in brackets). This observation indicates the necessity of performing multiple trials for comparing incomplete solvers.
BandMax MaxFPS NuWLS SATLike
MSE23-w 42 (30) 26 (20) 81 (81) 26 (18)
MSE23-uw 73 (42) 71 (34) 110 (99) 51 (30)
MSE22-w 49 (40) 49 (41) 103 (103) 35 (27)
MSE22-uw 89 (65) 86 (64) 106 (91) 55 (39)
Table 1: Number of instances in which a solver obtains the best performance among the competing solvers
Table 2 presents the scores of the tested solvers, of which values are averaged across the instances for each benchmark track. Since we conduct ten independent runs for each instance, the score of a solver for an instance $i$ is the average of ten trails. $score(i)=1$ indicates that a solver obtains the best solution for the instance $i$ among the competing solvers in all runs, and $score(i)=0$ indicates that no feasible solution is obtained. We do not take the scores of the instances in which no solver obtains a feasible solution into account. We observe that the orders of the solvers’ scores are identical to the ones in Table 1. In addition, according to Figure 1, there exist several instances in which no solver can obtain a feasible solution, and the solvers obtain the same or similar results in many instances, as shown in light yellow color, e.g., $score>0.9$. Therefore, the high variance in scores occurs in a limited number of instances. Still, Figure 1 presents that NuWLS obtains better scores in more instances compared to the other solvers.
BandMax MaxFPS NuWLS SATLike
MSE23-w 0.849 0.867 0.904 0.827
MSE23-uw 0.800 0.782 0.902 0.705
MSE22-w 0.869 0.889 0.942 0.864
MSE22-uw 0.854 0.852 0.901 0.731
Table 2: Aggregated scores of the solves
Figure 1: Heatmap illustrating the scores for individual instances of four benchmark tracks. Each column represents an instance, while each line represents a different solver. The color depicts the score achieved by the solvers, with lighter shades indicating better performance. The benchmark tracks “MSE23-w”, “MSE23-uw”, “MSE22-w” and “MSE22-uw” consist of 160, 179, 197, and 179 instances (as shown by columns) respectively. The names of the instances have been omitted from the y-axis due to the limited space.
3.2 Anytime Performance
The assessment in Section 3.1 provides initial insights into the performance of the tested solvers. However, these results only focus on the cutoff time of $300s$. While we can conduct similar analyses for multiple cutoff times, it remains challenging to aggregate and interpret the fixed-budget results across multiple cutoff time scenarios to understand the convergence of solvers. Therefore, we ask for an indicator that can assess anytime performance across various time budgets. Furthermore, this indicator shall be quantitative, allowing us to measure performance differences on a scale rather than by order. Particularly for instance-based scenarios such as MaxSAT, this indicator needs to be universal and independent of the scale of solution quality (e.g., cost of assignments) for each instance, enabling fair performance aggregation across instances.
Anytime Performance Metric.
We assess the anytime performance of MaxSAT solvers using the ECDF values of a set of cutoff times, where ECDF indicates the fraction of the obtained solutions satisfying a specific quality. We denote $\phi$ as the solution quality, e.g., the cost of an assignment. Given a set of solutions with $\Phi_{i}=\{\phi_{i1},\phi_{i2},\ldots\}$ for a problem instance $i$ and a solver $A$ that obtains the best-found solution with $\phi_{A_{i}}$ within the cutoff $t$, $A$’s ECDF value at $t$ is:
$\text{ECDF}(A,i,t)=\frac{\mid\{\phi\in\Phi_{i}\mid\phi\geq\phi_{A_{i}}\}\mid}{\mid\Phi_{i}\mid}$ (2)
We work on minimization in this paper, and $\{\phi\in\Phi_{i}\mid\phi\geq\phi_{A_{i}}\}$ denotes the subset of solutions that are not better than the best-found one of $A$ for instance $i$. For example, an algorithm $A$ obtains a set of pairs of optimization time and the corresponding best-found solution $O=\{(t_{1},\phi_{o1}),(t_{2},\phi_{o2}),\ldots\}$, which are plotted by blue dots in Figure 2. Given a set $\Phi$ of solutions plotted by red stars, we present on the right subfigure the number of instances in $\Phi$ that are not better than the corresponding $\phi_{o}$ for each $t$, and the values normalized by the size of $\Phi$ are ECDFs.
Figure 2: An example of ECDF calculation. Left: The blue dots represent pairs $(t,\phi)$, where a better solution with fitness $\phi$ ($y$-axis) is obtained at time $t$ ($x$-axis). Right: the ECDF values based on the given set of solution fitness as presented in red stars on the left.
Because ECDF is with a ratio scale, we can evaluate the ECDFs of a solver $A$ for a set of optimization times $t$ to measure $A$’s anytime performance. Additionally, ECDFs can be combined across multiple problem instances $i$ by considering specific $\Phi$ for each instance.
To calculate ECDFs, we form a set of solution fitness values $\Phi_{i}$ for each instance $i$. This solution fitness set comprises the costs of solutions that have been visited by any MaxSAT solver during our experiments. We compute the ECDFs at $100$ specific optimization times, chosen within the range of $[0,300s]$ following a logarithmic scale. This set of optimization times introduces a preference for evaluating the performance in terms of attaining high-quality solutions in less time.
Note that measuring ECDFs has been adopted in the black-box optimization community, where researchers quantify “time” in terms of function evaluations, i.e., the number of solutions that have been evaluated Hansen et al. (2016); Wang et al. (2022). Function evaluation ensures reproducibility and erases disturbances caused by the operation system, hardware environment, programming design, and other factors. However, MaxSAT solvers are usually applied in practical scenarios concerning cpu time, and cpu time has been commonly used when measuring MaxSAT solvers’ performance. We have conducted similar analyses by using cpu time and function evaluations as the time metric, and the results concerning both time metrics show consistent convergence progress by using the favored black-box benchmarking environment Wang et al. (2022); de Nobel et al. (2023). We provide corresponding data concerning function evaluations in our repository and present in this paper the results using cpu time as the time metric while ensuring that all experiments are conducted in identical environments.
(a) Left: WPMS (instances of MSE22-w and MSE23-w) and Right: PMS (instances of MSE22-uw and MSE23-uw)
(b) Left: “decision-tree” instances in MSE23-w ($15$ in total), and Right: “ParametricRBACMaintenance” instances in MSE23-w ($13$ in total).
Figure 3: Aggregated ECDFs of each solver for different sets of problem instances. $x$-axis indicates cpu time, and $y$-axis indicates the corresponding ECDF values.
Observing Anytime Performance.
Figure 2(a) shows that NuWLS outperforms the other solvers for both weighted and unweighted MaxSAT. BandMax and MaxFPS show similar performance, but by examining the “anytime” analysis we can see that MaxFPS slightly outperforms BandMax regarding within a time limit of $10s$, and the order of their performances alter as the optimization time increases for the weighted MaxSAT. Furthermore, detailed analysis of anytime performance for specific instances can reveal interesting findings. Although NuWLS demonstrates advantages over other solvers considering ECDFs aggregated across the entire set of tested problem instances, it does not exhibit superiority on $15$ “decision-tree” instances after being trapped in local optima around $10s$, as shown in Figure 2(b). On the other hand, while BandMax obtains similar performance to SATLike for $13$ “ParametricRBACMaintenance” instances when cpu time is less than $100s$, it outperforms the other solvers afterward. Compared to the fixed-budget performance assessment, these anytime performance observations offer more insights into the behavior of algorithms for specific problem instances or sets of instances. These observations can be valuable for enhancing algorithm design.
Figure 4: Heatmap illustrating the ECDFs for individual instances of four benchmark tracks. It follows the layout of Figure 1, but the color depicts the ECDF values achieved by the solvers.
BandMax MaxFPS NuWLS SATLike
MSE23-w 0.444 0.445 0.584 0.409
MSE23-uw 0.546 0.548 0.696 0.466
MSE22-w 0.458 0.470 0.624 0.420
MSE22-uw 0.589 0.602 0.696 0.494
Table 3: Aggregated ECDFs of the solvers for each benchmark track. The ECDF value for each run on an instance is normalized by the number of considered cutoff times, i.e., $100$. The presented values are average across the corresponding instances.
Quantitative Assessment with Variation
As mentioned previously, ECDF is with a ratio scale that allows aggregation of values across multiple optimization times to obtain quantitative assessments for each instance, which is the so-called (approximated) area under the ECDF curve (AUC) Ye et al. (2022). Table 3 presents the aggregated ECDFs for the four benchmark tracks, showing that the order of the solver’s assessment based on ECDF remains the same as the one based on scores, as presented in Table 2. When analyzing the behavior in specific instances, Figure 4 shows higher variances in ECDFs for each instance compared to the ones in scores (see Figure 1). This measn that instead of having a large set of scores higher than $0.9$, as spread across the yellow area of Figure 1, we can differentiate the solvers’ performance using ECDF for most instances, except those in which no feasible solution is obtained. Furthermore, high-variance ECDFs list across different instances for each solver. Moreover, the difference in such convergence process can be quantitatively estimated using ECDFs. Due to the space limit, we are unable to provide a table with complete ECDF values, but all the data is available in our repository.
4 Hyperparameter Optimization
Contemporary algorithms, such as heuristics, evolutionary algorithms, machine learning models, etc., are usually parametric. Therefore, parameter settings are essential for the competitive performance of algorithms. Before the development of automatic tools, parameter settings were manually chosen based on experts’ experience or repetitive testing using Grid Search. Nowadays, the tools such as SMAC Lindauer et al. (2022) and irace López-Ibáñez et al. (2016) have been commonly applied for automatic HPO. While practical scenarios may propose various requirements for HPO, we address in this paper the algorithm configuration (AC) problem, which meets the requirement of tuning MaxSAT solvers for a set(s) of problem instances. The Algorithm Configuration Lindauer et al. (2022) aims at searching a well-performance configuration, i.e., parameter setting, $\lambda\in\Lambda$ of an algorithm $A$ across a set of problem instances $\{i_{1},\ldots\}\subset\mathcal{I}$:
$\mathbf{\lambda}^{*}\in\underset{\lambda\in\Lambda}{\arg\min}\ e(\lambda)$ (3)
where the cost function $e$ is usually set as the best-found solutions’ quality intuitively when tuning MaxSAT LS solvers. However, in this section, we introduce ECDF as an alternative candidate for $e$ and compare the results obtained by using these two different mechanisms as the cost function for tuning MaxSAT LS solvers.
Settings of Hyperparameter Optimization.
We use the AC arcade of SMAC3  [^1] https://github.com/automl/SMAC3 to compare the results of tuning fixed-budget performance and anytime performance, namely, the quality of best-found solutions (Best-f) and ECDFs, respectively. Using Best-f as the cost function, it straightforwardly calculates the fitness of the best-found solution obtained within a given cutoff $t_{B}$. As for calculating ECDF, we form the a set $T$ of $50$ optimization times chosen within a range of $[0.1s,t_{B}]$ following a logarithmic scale. In Section 3, we prefer solvers to obtain high-quality solutions in less time when assessing anytime performance. However, the time budget $t_{B}$ for executing each run of configurations is usually relatively small for HPO scenarios, e.g., $t_{B}=100s<300s$ in this work. Therefore, following a bias toward the configurations that have the potential to yield better solutions when provided with an additional time budget, we use $t_{B}-T$ as the set of optimization times to calculate ECDFs. Regarding the target (i.e., solution quality) set $\Phi_{i}$ for each instance $i$, we start with a set $F_{i}$ of five fitness values chosen from $[f_{\text{init\_min}},f_{\text{init\_max}}]$ based on the linear scale, where $f_{\text{init\_min}}$ and $f_{\text{init\_max}}$ are the fitness of the best and the worst solutions visited by the initial configuration. When a configuration obtains a better best-found solution for an instance during the tuning process, the corresponding best-found fitness value will be added to $F_{i}$. In addition, the ECDF of a configuration for a given instance is calculated by aggregating the values across $50$ optimization times, without normalizing it by the size of the solution fitness set. In practice, we use a modified $\text{ECDF}^{\prime}=-\text{ECDF }\cdot\mid\Phi_{i}\mid$ measuring $e$ of configurations for each instance $i$. By using $\text{ECDF}^{\prime}$, it decreases accordingly when a better configuration achieves a better best-found solution, thereby increasing the size of $\Phi$. In this way, tuning Best-f and ECDF both are both framed as minimization tasks for HPO. Though computing ECDF is more complex than computing Best-f, in practice, this computing takes linear time, and its time consumption is negligible compared to the time budget allocated for each run of a configuration.
[…]
NuWLS SATLike
Instances Best-f ECDF Best-f ECDF
MSE23-w 0.386 0.450 0.139 0.237
0.398 0.455 0.184 0.246
0.409 0.456 0.325 0.329
0.421 0.465 0.328 0.410
0.457 0.503 0.329 0.412
Avg. 0.414 0.466 0.261 0.327
MSE23-uw 0.421 0.432 0.307 0.372
0.450 0.473 0.337 0.476
0.457 0.505 0.345 0.483
0.483 0.531 0.409 0.510
0.495 0.591 0.456 0.549
Avg. 0.461 0.506 0.371 0.478
[…]
(a) NuWLS
(b) SATLike
Figure 5: ECDFs of the NuWLS and SATLike configurations obtained by tuning for Best-f and ECDF for two sets of instances: Left: MSE23-w and Right: MSE23-uw. Results plotted in one line are from five configurations obtained by independent runs of SMAC.
NuWLS SATLike
Instances Best-f ECDF Best-f ECDF
MSE23-w 0.833 0.855 0.495 0.676
0.835 0.858 0.515 0.684
0.842 0.862 0.807 0.834
0.848 0.870 0.816 0.903
0.859 0.883 0.835 0.907
Avg. 0.843 0.866 0.694 0.801
MSE23-uw 0.814 0.823 0.566 0.654
0.818 0.837 0.577 0.830
0.829 0.845 0.584 0.837
0.840 0.856 0.733 0.844
0.846 0.891 0.757 0.860
Avg. 0.829 0.850 0.643 0.805
Table 5: Aggregated scores and the average of five NuWLS configurations and five SATLike configurations.
Anytime Performance Leads to Better Configurations.
We first compare the anytime performance of the obtained configurations. Table 4 presents the aggregated ECDFs across weighted and unweighted MaxSAT instances, and Figure 5 demonstrates the mean and $95\%$ confidence interval of ECDFs of five configurations along cpu time. Since we focus on the impact of the cost function, we analyze the comparison results separately for NuWLS and SATLike. These results are computed based on the corresponding ten configurations for each benchmark track. According to Table 4 and Figure 5, for both weighted and unweighted tracks, the optimal configuration among the ten ones is obtained by using ECDF as the cost function, and the results obtained by tuning for ECDF generally outperform the ones of tuning for Best-f. More interestingly, as observed in Table 5, tuning for ECDFs enables obtaining configurations that can achieve better scores than those obtained by tuning for the fitness of the best-found solutions. This suggests that using ECDF as the cost function can lead to better configurations in terms of both anytime performance and best-found solution quality.
One reason for this finding is that ECDFs can provide dense search space for HPO compared to Best-f, as shown in Figure 1 and Figure 4. For example, within a given time budget $t_{B}$, even though two different configurations $\lambda_{1}$ and $\lambda_{2}$ may obtain the same best-found solution for an instance $i$ using different time $t\leq t_{B}$, Best-f will deliver identical information to HPO tools, resulting in a sparse search space, but ECDF can distinguish these two configurations.
In addition, when conducting HPO, the time budget $t_{B}$ is usually set as a relatively small value due to time limits. For example, in our tuning process of SMAC, we allocated a budget of $100s$ for each configuration. However, when assessing the final obtained configurations, the budget is set as $300s$. In such scenarios where configurations are allocated with limited time budgets, via considering multiple cutoff times, ECDF can robustly estimate the potential ability of a configuration to achieve better solutions within a larger time budget. On the other hand, using Best-f as the cost function may be misled by fortunate results obtained at a specific cutoff time.
5 Conclusion
We have introduced ECDF for anytime performance analysis of MaxSAT local search solvers. Our assessments have been conducted for the four state-of-the-art solvers, namely BandMax, MaxFPS, NuWLS, and SATLike. We illustrate that ECDF can measure solvers’ anytime performance regarding multiple cutoff times, provide quantitative assessments with a ratio scale that can be aggregated across multiple instances, and differentiate the solvers that are considered identical in terms of fixed-budget performance. Although we have presented the application in the context of MaxSAT local search, we expect that the applied techniques can be transparent to the hybrid methods integrating complete and incomplete solvers. Our experiments show that ECDF can serve as a valuable metric for hyperparameter optimization to obtain better configurations of local search solvers. Additionally, by utilizing customized sets of time budgets and solution fitness to calculate ECDF, we anticipate that ECDF can serve as a valuable tool for researchers to comprehend the behavior of algorithms and improve their design for specific scenarios. Furthermore, we will study the algorithm portfolio configuration, which has a significant impact on MaxSAT, considering anytime performance of solvers.
6 Data Availability
The source code and raw data of the tested solvers, code and reproducible output of SMAC scenarios, post-processing results presented in the paper, the performance of tuning BandMax, and example data considering Function Evaluations as the time metric are publicly available in our repository: https://github.com/AcademicSubmission/IJCAI2024-p3562/.